OpenStudy (anonymous):

write the fraction 40/96 in simplest form

OpenStudy (anonymous):

what times what will get you 40.. they both have to be the same number though

OpenStudy (anonymous):

would it be 5/12

OpenStudy (anonymous):

oh good job yes it would!! great

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

you are welcome andy

OpenStudy (anonymous):

what about 14/16-6/16

OpenStudy (anonymous):

just susubtracting correct?

OpenStudy (anonymous):

yeah and in simplest form

OpenStudy (anonymous):

8/16 = 1/2 so 1/2 is simplest form

OpenStudy (anonymous):

oh ok thank you, what would you do if the denominator was different?

OpenStudy (anonymous):

like 5/7+3/8

OpenStudy (anonymous):

we'll take LCM

OpenStudy (anonymous):

7 * 8 = 56 so (5/7)*(8/8) (3/8)*(7/7) 40/56 + 21/56 now we have same denominator add them 61/56

OpenStudy (anonymous):

I hope you understand it :)

OpenStudy (anonymous):

ok just making sure its 3/8 not 8/8

OpenStudy (anonymous):

i kinda dont get the second part parter you got 56

OpenStudy (anonymous):

after you got 56

OpenStudy (anonymous):

oh its simple I just multiplied the whole fraction with 7 or 8 to make the denominators 56

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

yes ok thanks so much!

OpenStudy (anonymous):

\[\frac{ 5 }{ 7}\times \frac{ 8 }{ 8 } + \frac{ 3 }{ 8}\times \frac{ 7 }{ 7 }\]

OpenStudy (anonymous):

so that you can understand better :D welcome